If lithium ion Li2+ is treated as a classical system of an electron of charge ?e going around a nucleus of charge 3e on a circle of radius 1.77
Answer #1

The magnetic moment is defined by:

\mu = I. A

for the electron

I= \frac{e}{T}= \frac{e}{2\pi/\omega} = \frac{e\omega}{2\pi}

A = \pi r^2

we get:

\mu = \frac{e\omega r^2}{2} (1)

using the classical system:

F = \frac{ke(3e)}{r^2} = ma_c = m\omega^2 r

we get:

\omega = \sqrt{\frac{3ke^2}{mr^3}}

using this expression in

\mu = \frac{e r^2\sqrt{\frac{3ke^2}{mr^3}}}{2} = \frac{e^2}{2}\sqrt{\frac{3kr}{m}}

using values we get:

\mu =9.28 \times 10^{-24} Am^2